OpenStudy (lena772):

3x+2y=15 Determine any point that is a solution and will be on this line. Justify your answer algebraically.

ganeshie8 (ganeshie8):

set any value to x, and find the corresponding y value

OpenStudy (lena772):

3x+2y=15x?

ganeshie8 (ganeshie8):

for example :- set x = 0 3*0 + 2y = 15 2y = 15 y = 15/2 = 7.5

ganeshie8 (ganeshie8):

so the point (0, 7.5) is a solution to the given equation, and lies on the line as well

OpenStudy (lena772):

Thanks !